# ICCO: Learning an Instruction-conditioned Coordinator for Language-guided Task-aligned Multi-robot Control

### v1

@@ IV-B Policy model
Furthermore, the local agent has one local policy that computes the action by
$a^{i}_{t}=\operatorname{argmax}_{a^{i}_{t}}Q^{i}(\tau^{i}_{t},a_{t}^{i}|z^{i}_{t}),$ (5)
where the action $a^{i}_{t}$ is selected to maximize the local action-value function $Q^{i}$ based on the local observation $o^{i}_{t}$ and individual TACI $z^{i}_{t}$. Following QMIX, this function is modeled by the joint action-value function $Q^{\rm tot}_{\theta}$ as in Eq. (1).
@@ IV-C Training objectives
where $\bm{\zeta}_{t}=\left(\bm{o}_{t+1},\bm{a}_{t+1},\ldots,\bm{o}_{t+T},\bm{a}_{t+T}\right)$ represents the observation-action trajectory from step $t+1$ to $t+T$.
@@ IV-D Training without LLM and execution with LLM
To address this issue, we adopt a training approach where policies are learned without LLMs during training but then executed with LLMs, which are adopted only during this latter phase following the method proposed previously [16]. Specifically, during the training phase, policies are trained by random sampling within the expected instruction vector space without using an LLM. During the execution phase, the LLM converts diverse language instructions into instruction vectors, which are then used to execute the policy. This approach reduces training time, compared to methods that incorporate LLMs during training, while still maintaining the capability to handle various language instructions.
@@ V-A1 Simulation Setup
Each agent can observe only other agents, resources, the invader and the home within a circle of radius 0.65 m. The instruction vector $\bm{v}_{t}$ consists of the instruction vectors for all agents, where each agent’s instruction is specified by waypoints. The agent’s reward is given by $r=r_{\rm task}+r_{\rm inst}$, where $r_{\rm task}$ and $r_{\rm inst}$ represent the task and instruction-following rewards, respectively.
The task reward is given by $r_{\rm task}=r_{\rm pick}+r_{\rm collect}+r_{\rm defense}$, where $r_{\rm pick}=5$ for picking a resource, $r_{\rm collect}=1$ for releasing it at home, $r_{\rm defense}=4$ for hitting an invader, and $r_{\rm defense}=-4$ if the invader reaches home. The instruction-following reward is given by $r_{\rm inst}=1.3(e_{\mathrm{cossim}}+0.1e_{\mathrm{dist}})$, where $e_{\mathrm{cossim}}$ is the cosine similarity between two displacement vectors: one from the agent’s current position to the next position and one from the nearest waypoint to the second-nearest waypoint, and $e_{\mathrm{dist}}$ is the distance between the agent’s current position and the nearest waypoint.
@@ V-D Can the Coordinator in ICCO be replaced by an LLM with task-aligned prompt engineering?
Setup: In this experiment, we further investigate the effectiveness of ICCO by comparing it to an approach that employs task-aligned prompts as a complementary method instead of using a Coordinator. Since the previous two experiments confirmed instruction following, this experiment focuses on evaluating task performance. We compared our method with the following methods:
@@
These methods commonly utilized a task-aligned prompt that includes resources, invaders, and task rewards. See the Appendix for further details on the prompt.
@@
Overall, the Coordinator in ICCO cannot be replaced by an LLM with task-aligned prompt engineering.
TABLE II: Comparison of the number of picks, collections, and invader defenses for 20 trials
@@ VII CONCLUSION
This paper proposed ICCO, a MARL framework that balances language-instruction following and cooperative task execution. ICCO consists of a Coordinator, which generates TACI from language inputs and global observations. The coordination and local policies are consistently trained within the CTDE framework to balance instruction following and task requirements.
Experiments show that ICCO outperforms baseline methods in balancing instruction following and task execution. It also effectively handles natural language instructions with an LLM. Notably, the Coordinator cannot be replaced by an LLM with prompt engineering, highlighting the limitations of relying solely on prompts for task execution.
As future work, we aim to develop a general approach for converting natural language into numerical representations and to evaluate this approach across a broader range of tasks.

### v2

@@ IV-B Policy model
Furthermore, the local agent has one local policy that computes the action by $a^{i}_{t}=\operatorname{argmax}_{a^{i}_{t}}Q^{i}(\tau^{i}_{t},a_{t}^{i}|z^{i}_{t}),$ where the action $a^{i}_{t}$ is selected to maximize the local action-value function $Q^{i}$ based on the local observation $o^{i}_{t}$ and individual TACI $z^{i}_{t}$. Following QMIX, this function is modeled by the joint action-value function $Q^{\rm tot}_{\theta}$ as in Eq. (1).
@@ IV-C Training objectives
where $\bm{\zeta}_{t}=\left(\bm{o}_{t+1},\bm{a}_{t+1},\ldots,\bm{o}_{t+T-1},\bm{a}_{t+T-1}\right)$ represents the observation-action trajectory from step $t+1$ to $t+T-1$.
@@ IV-D Training without LLM and execution with LLM
To address this issue, we adopt a training approach where policies are learned without LLMs during training but then executed with LLMs, which are adopted only during this latter phase following the method proposed previously [16]. Specifically, during the training phase, policies are trained by random sampling within the expected instruction vector space without using an LLM. Each instruction vector is generated as a smooth trajectory by sampling Gaussian noise $\bm{\epsilon}_{k}$ at each time step $k$ and computing the cumulative sum from the agent’s current position, given by $\bm{v}_{t+k}=\bm{p}_{t}+\sum^{K}_{k=1}\bm{\epsilon}_{k}$. During the execution phase, the LLM converts diverse language instructions into instruction vectors for each agent. These vectors are clipped within a predefined value to ensure that agents stay within the movable region, and to prevent the LLM outputs from becoming out-of-distribution relative to the training data. This approach reduces training time, compared to methods that incorporate LLMs during training, while still maintaining the capability to handle various language instructions.
@@ V-A1 Simulation Setup
Each agent can observe only other agents, resources, the invader and the home within a circle of radius 0.65 m. The instruction vector $\bm{v}_{t}$ consists of the instruction vectors for all agents, where each agent’s instruction is specified by waypoints. The global state provided to the coordination policy includes the positions and velocities of all agents, binary flags indicating whether they are carrying resources or defending invaders, the positions of environmental elements such as resources and invaders, and instruction vectors generated by the LLM for all agents.
The agent’s reward is given by $r=r_{\rm task}+r_{\rm inst}$, where $r_{\rm task}$ and $r_{\rm inst}$ represent the task and instruction-following rewards, respectively. The task reward is given by $r_{\rm task}=r_{\rm pick}+r_{\rm collect}+r_{\rm defense}$, where $r_{\rm pick}=5$ for picking a resource, $r_{\rm collect}=1$ for releasing it at home, $r_{\rm defense}=4$ for hitting an invader, and $r_{\rm defense}=-4$ if the invader reaches home. The instruction-following reward is given by $r_{\rm inst}=1.3(e_{\mathrm{cossim}}+0.1e_{\mathrm{dist}})$, where $e_{\mathrm{cossim}}$ is the cosine similarity between two displacement vectors: one from the agent’s current position to the next position and one from the nearest waypoint to the second-nearest waypoint, and $e_{\mathrm{dist}}$ is the distance between the agent’s current position and the nearest waypoint.
@@ V-D Can the Coordinator in ICCO be replaced by an LLM with task-aligned prompt engineering?
Setup: In this experiment, we further investigated the effectiveness of ICCO by comparing it to an approach that employs task-aligned prompts as a complementary method instead of using a Coordinator. Since the previous two experiments confirmed instruction following, this experiment focused on evaluating task performance. We compared our method with the following methods:
@@
These methods commonly used a task-aligned prompt that includes resources, invaders, and task rewards. See Appendix for further details on the prompt.
@@
In summary, the Coordinator in ICCO cannot be replaced by an LLM with task-aligned prompt engineering.
TABLE II: Comparison of the numbers of picks, collections, and invader defenses over 20 trials
@@ VII CONCLUSION
This paper proposed ICCO, a MARL framework that balances language-instruction following and cooperative task execution. ICCO consists of a Coordinator, which generates TACI from language inputs and global observations. The coordination and local policies are consistently trained within the CTDE framework to balance instruction following and task requirements. Experiments show that ICCO outperforms baseline methods in balancing instruction following and task execution. It also effectively handles natural language instructions with an LLM. Notably, the Coordinator cannot be replaced by an LLM with prompt engineering, highlighting the limitations of relying solely on prompts for task execution. As future work, we aim to develop a general approach for converting natural language into numerical representations and to evaluate this approach on more challenging tasks, including cooperative manipulation.